OpenStudy (anonymous):

\[Solve: \int\limits_{}^{}3\div(2x^2-x-1)dx\] This is driving me nuts! :(

OpenStudy (turingtest):

either try to factor it and use partial fractions, or complete the square then trig sub

OpenStudy (anonymous):

I get the answer -2*ln(2x+1)+ln(x-1)+c Supposedly, the correct answer is ln(x-1)-ln(2x+1)+c

OpenStudy (anonymous):

I went the way of partial fractions.

OpenStudy (turingtest):

look like you used partial fractions...

OpenStudy (turingtest):

...incorrectly

OpenStudy (experimentx):

couldn't think of factoring ... looks like factoring.

OpenStudy (anonymous):

Haha, trouble is; I can't find where I'm wrong :(

OpenStudy (experimentx):

(2x+1)(x-1)

OpenStudy (anonymous):

Indeed.

OpenStudy (turingtest):

so let's see what you did in the PF part

OpenStudy (turingtest):

\[{A\over2x+1}+{B\over x-1}={3\over(2x+1)(x-1)}\implies A(x-1)+B(2x+1)=3\]

OpenStudy (turingtest):

so easiest is probably to choose some values for x... which ones did you choose?

OpenStudy (anonymous):

a+2b=0 b-a=3

OpenStudy (turingtest):

that's the other way of doing it which is why you probably messed up try plugging in x=1 into my post above

OpenStudy (turingtest):

either way you get the same answer

OpenStudy (anonymous):

The book I'm using only explained it one way. :/

OpenStudy (turingtest):

I'm explaining another :)

OpenStudy (anonymous):

I do get B=1 and then a=-2

OpenStudy (turingtest):

yes you do

OpenStudy (anonymous):

Which then - leads to: INT(-2/(2x+1))+INT(1/(x-1))

OpenStudy (anonymous):

dx

OpenStudy (turingtest):

oh wait you have the right setup the answer is right

OpenStudy (anonymous):

But according to both wolf and the book, the other answer is correct, are they the same? This is crazy! :D

OpenStudy (turingtest):

you forgot the u-sub on the first term is all :)

OpenStudy (turingtest):

I almost missed it myself for a moment....

OpenStudy (anonymous):

Could you elabourate on the u-sub part? I don't follow.

OpenStudy (turingtest):

myin will I'm sure

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{2x+1} dx => u=2x+1 => du=2 dx\]

OpenStudy (anonymous):

Ah! Thanks so much guys. :)

OpenStudy (turingtest):

details, details... the death of mathematicians everywere

OpenStudy (anonymous):

Haha, partial fractal decomposition is so far the most tedious I've done. Takes sooooo long and easy to mess up.

OpenStudy (turingtest):

did you understand the method I was describing above? that way is often easier than solving a system, as you were taught

myininaya (myininaya):

I think this was is easier than trig sub Do you know trig sub?

OpenStudy (turingtest):

I meant the other way to find the coefficients

myininaya (myininaya):

I was responding to noliec's comment on partial fractions being most tedious

OpenStudy (turingtest):

oh ic

OpenStudy (anonymous):

I think so TuringTest! :) As for myininaya, I don't really know what you mean. :D

myininaya (myininaya):

But in response to you i'm addicted to solving the system lol

myininaya (myininaya):

Ok well I'm sure you will figure out in this class This is like cal 2 right?

OpenStudy (anonymous):

Trigonometric substitution? I've done some, not all too much though. I'm a Swedish student at senior high school.

OpenStudy (anonymous):

Hence the trouble converting from "Swedish typestyle" math to English, hehe.

OpenStudy (anonymous):

I would do this with http://en.wikipedia.org/wiki/Heaviside_cover-up_method

OpenStudy (anonymous):

I will probably go into details of alternative methods in this chapter of the book, we shall see! :)